# Inference of Dependency Knowledge Graph for Electronic Health Records

### v1

@@ 3 Theoretical Properties
Assumption 3.1 (Scaling).
Assume that there exist constants $0<c<C$ such that $p\geq c\log^{2}d$ and $\log p\leq C\log(\log d)$. We suppose $T_{i}=T$ for each $i$, and we have $cnd\log^{2}d\leq T\leq Cd^{4\log d}$ and $\log d/\log n=o(1)$.
@@
Assumption 3.1 states the necessary scaling conditions to guarantee that the estimator converges to the truth at a fast enough rate. The lower bound for $p$ is to control the potential information loss caused by low-rank approximation, and the upper bound for $p$ is to control the variance of the estimator $\widetilde{\mathbb{PMI}}$ and maintain the “near low-rank” structure of the population PMI matrix. Although we could consider $\{T_{i}\}_{i=1}^{n}$ to be of the same order, leading to $T_{i}/T_{j}=O(1)$ for any $i,j\in[n]$, we opt for the simpler assumption of $T_{i}=T$ for the ease of expression. Here we impose the upper bound for $T$ because when $T$ grows exceptionally large, the advantages of employing the low-rank estimator $\widetilde{\mathbb{PMI}}$ over its empirical counterpart $\widehat{\mathbb{PMI}}$ disappear. The intricate Markov dependency structure in (2.2) necessitates the scaling assumption on the sample size $n$, ensuring the signal is significant enough to conduct the inference.
@@
The upper bound for $\|\mathbf{U}\|_{2,\infty}$ imposes the incoherence condition (Candès and Tao, 2010) on the eigen-space of $\mathbf{V}$. Concurrently, the lower bound for $\min_{i}\|\mathbf{U}_{i,\cdot}\|$ is very weak since it is of order $p^{2}/\xi d$, which decays much faster than the upper bound combined with Assumption 3.3. These two inequalities together guarantee that the eigenvectors are roughly uniformly distributed in the eigen-space.
@@
Suppose Assumptions 3.1-3.3 hold, then we have the entry-wise asymptotic normality of the residual matrix $\mathbf{W}$, i.e., for any entry $(w,w^{\prime})\in[d]\times[d]$, we have
$\frac{\sqrt{n}}{\sigma_{w,w^{\prime}}}\mathbf{W}_{w,w^{\prime}}\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(0,1)\text{ as }n,d,T\rightarrow\infty,$ (3.1)
where $\sigma_{w,w^{\prime}}^{2}=\text{Var}\big{(}\mathbb{C}_{w,w^{\prime}}^{(i)}/{\mathbb{E}}[\mathbb{C}_{w,w^{\prime}}^{(i)}]-\mathbb{C}_{w}^{(i)}/{\mathbb{E}}[\mathbb{C}_{w}^{(i)}]-\mathbb{C}_{w^{\prime}}^{(i)}/{\mathbb{E}}[\mathbb{C}_{w^{\prime}}^{(i)}]\big{)}$ is of order $d^{2}/T$.
@@
Let $\widetilde{\bm{\Sigma}}=(\mathbf{P}^{\star}\circ\mathbf{P}^{\star})\bm{\Sigma}_{E}+\bm{\Sigma}_{E}(\mathbf{P}^{\star}\circ\mathbf{P}^{\star})$. Under Assumptions 3.1-3.3, for the low-rank estimator $\widetilde{\mathbb{PMI}}$, we have for any entry $(w,w^{\prime})\in[d]\times[d],$
$\frac{\sqrt{n}}{\widetilde{\bm{\Sigma}}^{1/2}_{w,w^{\prime}}}(\widetilde{\mathbb{PMI}}_{w,w^{\prime}}-\alpha_{p}\mathbf{V}_{w}^{\top}\mathbf{V}_{w^{\prime}})\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(0,1)\text{ as }n,d,T\rightarrow\infty,$
where $\|\widetilde{\bm{\Sigma}}\|_{\max}=O(dp/(nT))$.
The proof of Theorem 3.6 can be found in Supplementary S2.3. In fact, the proof of Theorem 3.6 provides a Bahadur decomposition of $\widetilde{\mathbb{PMI}}-\alpha_{p}\mathbf{V}_{w}^{\top}\mathbf{V}_{w^{\prime}}$ to establish the entrywise asymptotic normality for the low-rank estimator $\widetilde{\mathbb{PMI}}$, which can be generalized for multiple testing. Compared with the variance of the empirical estimator $\sigma_{w,w^{\prime}}^{2}=d^{2}/T$ shown in Lemma 3.5, the variance of the low-rank estimator $\widetilde{\bm{\Sigma}}_{w,w^{\prime}}=O(dp/(nT))$ is much smaller. This suggests the superior performance of the low-rank estimator over the empirical estimator.
@@ 4.1 Numerical Performance of PMI Estimators
    Estimate the marginal occurrence probabilities $\bm{p}$ using the Monte Carlo (MC) method with $10^{7}$ i.i.d MC samples. Denote it as $\bm{p}_{\text{mc}}$.
@@
Here the last step is to guarantee that $\mathbf{V}^{\top}\bm{p}=0$. Then we generate $100$ i.i.d replicates. For each replicate, the sequence $\{w_{i,t}\}_{1\leq i\leq n,1\leq t\leq T}$ is generated in accordance with (2.1), and $\|\widehat{\mathbb{PMI}}-\alpha_{p}\mathbf{V}\mathbf{V}^{\top}\|_{\max}$, $\|\widetilde{\mathbb{PMI}}-\alpha_{p}\mathbf{V}\mathbf{V}^{\top}\|_{\max}$, and $\|\mathbb{PMI}-\alpha_{p}\mathbf{V}\mathbf{V}^{\top}\|_{\max}$ are calculated. As evidenced by the first plot in Figure 1, the low-rank estimator has less variability and outperforms the empirical estimator, and the bias $\|\mathbb{PMI}-\alpha_{p}\mathbf{V}\mathbf{V}^{\top}\|_{\max}$ is negligible compared to $\|\widehat{\mathbb{PMI}}-\alpha_{p}\mathbf{V}\mathbf{V}^{\top}\|_{\max}$ and $\|\widetilde{\mathbb{PMI}}-\alpha_{p}\mathbf{V}\mathbf{V}^{\top}\|_{\max}$, which are consistent with our theoretical results.
@@
Figure 1: Left plot is the median and the $20\%$-$80\%$ quantile range of max norm distances between word embeddings and PMI estimators across 100 independent runs. Middle and right plots are Q-Q plots of one standardized entry for the empirical (green) estimator and the low-rank (blue) estimator. The Middle: $(d,n,T,\kappa)=(100,1000,800,1),(w,w^{\prime})=(1,1)$. Right: $(d,n,T,\kappa)=(100,1000,1200,1),(w,w^{\prime})=(1,1)$.
@@ 4.2 Numerical Results on Hypothesis Testing
In this section, we simulate data following the global null distribution given in Section 2.3.2. For simplicity, we assume that each feature occurs with equal probability, such that $p_{w}=1/d$ for $1\leq w\leq d$. The sets $\{w_{i,t}\}_{1\leq i\leq n,1\leq t\leq T}$ are independently and identically drawn from a $\text{Multinomial}(1,(1/d,\cdots,1/d))$ distribution, considering $d\in\{100,200,300\}$ and $n\in\{200,400,800,1600\}$. Under these conditions, it is not possible to define the embedding, and thus our analysis is confined to the empirical estimator. We apply the methods described in Sections 2.3.1 and 2.3.2 to estimate the variance of each entry in $\widehat{\mathbb{PMI}}$, employing the Monte-Carlo technique to estimate the empirical type I error. This involves generating $n$ i.i.d longitudinal code sequences $100$ times, and subsequently calculating the type I errors for each method. When patient-level data is accessible, we estimate the variance of each entry using (2.9). In contrast, when patient-level data is unavailable, we estimate the variance of each entry via (2.11).
The testing is conducted on the ten edges from the $1$st entity to the $2$nd, $3$rd, $\ldots$, and $11$th entities, with the nominal significance level set as $\alpha=0.05$ for each edge. We then average these type I errors to estimate the empirical significance level for any arbitrary edge test, with the results presented in Table 1. It is observed that the averages of the ten empirical type I errors, as calculated by both methods, tend to approach $0.05$ when $n$ becomes sufficiently large. This observation underscores the asymptotic normality of the empirical estimator and suggests that even in the absence of patient-level data, it is feasible to estimate the variance of each edge and establish confidence intervals for individual edges.
$d$ 100 200 300
$n$ 200 400 800 200 400 800 200 400 800 1600
Patient level 0.072 0.061 0.042 0.069 0.067 0.049 0.077 0.077 0.064 0.040
Summary stat. 0.062 0.064 0.048 0.068 0.054 0.050 0.077 0.069 0.062 0.045
Table 1: The averages of the first ten edges’ empirical type I errors under the global null distribution using the empirical estimator. The row of “Patient level” uses the patient-level data in Section 2.3.1 and the row of “ Summary stat.” applies the method in Section 2.3.2 using the total co-occurrence matrix to estimate the variances.
@@ 4.3 Numerical Results on the Asymptotic Normality
In this section, we generate data from (2.1). The embeddings are constructed using $\mathbf{V}=d^{-\kappa}\mathbf{U}$, where $\kappa\in\{1,2\}$ and $\mathbf{U}$ is formed from the orthonormal bases of a $d\times p$ matrix with each element independently drawn from a standard normal distribution.
@@ 5 Applications to Electronic Health Records
We validate the efficacy of our algorithm using EHR data from the Veterans Affairs (VA) Healthcare System consisting of $12.6$ million patients who had at least one visit between $2000$ and $2019$. The dataset includes four specific domains of codified information: PheCodes for diseases, Current Procedural Terminology (CPT) codes for procedures, lab test codes, and RxNorm codes for medication prescriptions. Codes occurring fewer than $5000$ times were excluded to minimize noise from infrequent codes, resulting in a final count of $1776$ PheCodes, $224$ CCS codes, $6025$ lab codes, and $1561$ RxNorm codes as detailed in Hong et al. (2021). These codes were then utilized to construct a co-occurrence matrix, adopting a window size, $q$, of $30$ days following the analysis of Hong et al. (2021).
@@
From the results presented in Table 4, it is evident that our algorithm demonstrates a high power in detecting known-relation pairs. In particular, for several pair categories, such as PheCode-PheCode pairs from PheCode Hierarchy, and related disease-disease pairs, our algorithm achieves an exceptional power of over $85\%$ while all other methods have power lower than $75\%$. Furthermore, our method consistently outperforms the selection based on empirical PMI or cosine-similarity from other benchmarks for all pair categories. This underscores the capability of our algorithm to not only establish a threshold for selecting significant edges but also to leverage the variance information of the estimator, surpassing the performance of methods reliant solely on cosine similarity. Similar patterns can be found in Table LABEL:tab:app_power_pairs in Supplementary S3 with target FDR set as $0.1$.
@@
[…]
@@
To validate the accuracy of the KG, we conduct a manual review of the top twenty entities with the smallest $p$-values. This group includes eight PheCodes (diseases) and twelve RXNORMs (drugs). These entities are enumerated in Table LABEL:tab:AD_test_app in Supplementary S3. A substantial body of literature (Sosa-Ortiz et al., 2012; Bathgate et al., 2001; Nussbaum and Ellis, 2003; Reisberg et al., 2003; Tsuno, 2009; Olin and Schneider, 2002; Xu et al., 2008; Rabinowitz et al., 2007) confirms that these entities have a strong association with AD, thereby validating the precision of our KG.
To further assess the effectiveness of the proposed algorithm in identifying entities associated with the given disease AD, we incorporate a comparison with advanced language models ChatGPT (OpenAI, 2023a) and GPT-4 (OpenAI, 2023b). Specifically, we first select the top $50$ entities related to AD based on different methods, and then randomly sample an equal number of entities from the remaining set as negative samples. We then prompt ChatGPT and GPT-4 to rate the clinical relatedness between these $100$ entities and the target disease AD, outputting a score on a scale of 00 to $1$ for each entity. The evaluation involves computing the Spearman rank correlation between the negative $p$-values generated by our algorithm and the scores from ChatGPT/GPT4. Additionally, we examine correlations between cosine similarities obtained from alternative benchmarks and the ChatGPT/GPT scores, as detailed in Table 5. The results indicate that the rank correlation coefficients for KNIT with ChatGPT and GPT-4 are $0.553$ and $0.582$, respectively, outperforming other methods which all showed correlations less than $0.45$ with the language models. For reference, the rank correlation between ChatGPT and GPT-4 themselves is $0.610$.

### v2

@@ 3 Theoretical Properties
Assumption 3.1 (Scaling Conditions).
There exist constants $0<c<C$ such that $p\geq c\log^{2}d$ and $\log p\leq C\log(\log d)$. Additionally, we assume that $T_{i}=\Omega(p^{2}\log(d))$ for $i\in[n]$.
@@
Assumption 3.1 states the necessary scaling conditions to guarantee that the estimator converges to the truth at a fast enough rate. The lower bound for $p$ is to control the potential information loss caused by low-rank approximation, and the upper bound for $p$ is to control the variance of the estimator $\widetilde{\mathbb{PMI}}$ and maintain the “near low-rank” structure of the population PMI matrix. The intricate Markov dependency structure in (2.2) necessitates the lower bound assumption on the sequence length $T_{i},1\leq i\leq n$, ensuring the mixing property of probability density functions of the $p$-dimensional Markov chain $\{{\boldsymbol{c}}_{i,t}\}_{t=1}^{T_{i}}$.
@@
The upper bound for $\|\mathbf{U}\|_{2,\infty}$ imposes the incoherence condition (Candès and Tao, 2010) on the eigen-space of $\mathbf{V}$. Concurrently, the lower bound for $\min_{i}\|\mathbf{U}_{i,\cdot}\|$ is weak since it is of order $p^{2}/\xi d$, which decays much faster than the upper bound combined with Assumption 3.3. These two inequalities together guarantee that the eigenvectors are roughly uniformly distributed in the eigen-space.
@@
Suppose Assumptions 3.1-3.3 hold and $nT\geq T_{i}d^{2}\log^{2}d$ for all $i\in[n]$, then we have the entry-wise asymptotic normality of the residual matrix $\mathbf{W}$, i.e., for any entry $(w,w^{\prime})\in[d]\times[d]$, we have
$\frac{\sqrt{n}}{\sigma_{w,w^{\prime}}}\mathbf{W}_{w,w^{\prime}}\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(0,1),$ (3.1)
where
$\sigma_{w,w^{\prime}}^{2}=\frac{1}{n}\sum_{i=1}^{n}(\frac{T_{i}-q}{T-q})^{2}\text{Var}\big(\mathbb{C}_{w,w^{\prime}}^{(i)}/{\mathbb{E}}[\mathbb{C}_{w,w^{\prime}}^{(i)}]-\mathbb{C}_{w}^{(i)}/{\mathbb{E}}[\mathbb{C}_{w}^{(i)}]-\mathbb{C}_{w^{\prime}}^{(i)}/{\mathbb{E}}[\mathbb{C}_{w^{\prime}}^{(i)}]\big)$
is of order $d^{2}/T$. In the special case where $T_{i}=T$ for all $i\in[n]$, the asymptotic normality in (3.1) holds under Assumptions 3.1-3.3 without requiring a lower bound on the sample size $n$. In other words, $n$ does not necessarily need to approach infinity for the result to hold.
@@
Let $\widetilde{\bm{\Sigma}}=(\mathbf{P}^{\star}\circ\mathbf{P}^{\star})\bm{\Sigma}_{E}+\bm{\Sigma}_{E}(\mathbf{P}^{\star}\circ\mathbf{P}^{\star})$. Under Assumptions 3.1-3.3 and $nT\geq T_{i}d^{2}\log^{2}d$ for all $i\in[n]$, for the low-rank estimator $\widetilde{\mathbb{PMI}}$, we have for any entry $(w,w^{\prime})\in[d]\times[d],$
$\frac{\sqrt{n}}{\widetilde{\bm{\Sigma}}^{1/2}_{w,w^{\prime}}}(\widetilde{\mathbb{PMI}}_{w,w^{\prime}}-\alpha_{p}\mathbf{V}_{w}^{\top}\mathbf{V}_{w^{\prime}})\stackrel{{\scriptstyle d}}{{\longrightarrow}}N(0,1)\text{ as }n,T,d\rightarrow\infty,$
where $\|\widetilde{\bm{\Sigma}}\|_{\max}=O(\frac{dp^{2}}{T\xi^{4}})$.
Remark 3.5.
Although both $n$ and $T$ need to go to infinity, the growth requirement for $T$ is significantly weaker compared to $n$. The key constraint arises from the bias term $\mathbb{PMI}-\alpha_{p}\mathbf{V}\mathbf{V}^{\top}$, which depends only on $d$. Additionally, since the mixing properties of the Markov chain $\{p_{i,w}(t)\}_{t=1}^{T_{i}}$ depend on $d$, the lower bound $T_{i}=\Omega(p^{2}\log(d))$ ensures that $T_{i}$ must increase with $d$. However, this increase occurs at a much slower rate compared to $n$, which scales as $\Omega(d^{2}\log^{2}(d))$ when $T_{i}\approx T$. This distinction highlights that while $T$ grows with $d$, its growth is far outpaced by that of $n$, aligning with practical data scenarios.
In particular, the EHR dataset analyzed in Section 5 consists of $120,518$ patients, with sequence length quantiles of $370$, $1090$, and $2642$ at the $25$th, $50$th, and $75$th percentiles, respectively, and an average sequence length $T=2197.7$. The vocabulary size $d$ is $9,586$. These statistics confirm that the sample size is much larger than the vocabulary size and both the average and typical sequence lengths, supporting the practical validity of our assumptions.
The proof of Theorem 3.6 can be found in Supplementary S2.3. In fact, the proof of Theorem 3.6 provides a Bahadur decomposition of $\widetilde{\mathbb{PMI}}-\alpha_{p}\mathbf{V}_{w}^{\top}\mathbf{V}_{w^{\prime}}$ to establish the entrywise asymptotic normality for the low-rank estimator $\widetilde{\mathbb{PMI}}$, which can be generalized for multiple testing. Compared with the variance of the empirical estimator $\sigma_{w,w^{\prime}}^{2}=d^{2}/T$ shown in Lemma 3.5, the variance of the low-rank estimator $\widetilde{\bm{\Sigma}}_{w,w^{\prime}}=O(dp^{2}/(T\xi^{4}))$ is much smaller. This suggests the superior performance of the low-rank estimator over the empirical estimator.
@@ 4.1 Numerical Performance of PMI Estimators
    Estimate the marginal occurrence probabilities $\bm{p}$ using the Monte-Carlo (MC) method with $10^{7}$ i.i.d MC samples. Denote it as $\bm{p}_{\text{mc}}$.
@@
Here the last step is to guarantee that $\mathbf{V}^{\top}\bm{p}=0$. Then we generate $100$ i.i.d replicates. For each replicate, the sequence $\{w_{i,t}\}_{1\leq i\leq n,1\leq t\leq T}$ is generated in accordance with (2.1), and $\|\widehat{\mathbb{PMI}}-\alpha_{p}\mathbf{V}\mathbf{V}^{\top}\|_{\max}$, $\|\widetilde{\mathbb{PMI}}-\alpha_{p}\mathbf{V}\mathbf{V}^{\top}\|_{\max}$, and $\|\mathbb{PMI}-\alpha_{p}\mathbf{V}\mathbf{V}^{\top}\|_{\max}$ are calculated.
As evidenced by the first plot in Figure 1, the low-rank estimator has less variability and outperforms the empirical estimator, and the bias $\|\mathbb{PMI}-\alpha_{p}\mathbf{V}\mathbf{V}^{\top}\|_{\max}$ is negligible compared to $\|\widehat{\mathbb{PMI}}-\alpha_{p}\mathbf{V}\mathbf{V}^{\top}\|_{\max}$ and $\|\widetilde{\mathbb{PMI}}-\alpha_{p}\mathbf{V}\mathbf{V}^{\top}\|_{\max}$, which are consistent with our theoretical results.
@@
Figure 1: Left plot is the median and the $20\%$-$80\%$ quantile range of max norm distances between EHR entity embeddings and PMI estimators across 100 independent runs. Middle and right plots are Q-Q plots of one standardized entry for the empirical (green) estimator and the low-rank (blue) estimator. The Middle: $(d,n,T,\kappa)=(100,1000,800,1),(w,w^{\prime})=(1,1)$. Right: $(d,n,T,\kappa)=(100,1000,1200,1),(w,w^{\prime})=(1,1)$.
To evaluate the robustness of our method under model mis-specification, we conducted additional experiments where the original AR(1) process for the discourse vectors was replaced with alternative processes. Specifically, we considered a random walk on a unit sphere, following the approach in Lu et al. (2023), and an ARMA(1,3) process: ${\boldsymbol{c}}_{t+1}=\sqrt{\alpha}{\boldsymbol{c}}_{t}+\sqrt{1-\alpha}\bm{r}_{t+1}+0.2\bm{r}_{t}+0.1\bm{r}_{t-1}+0.05\bm{r}_{t-2}.$ The results, presented in Figure 2, indicate that these changes have minimal impact on PMI estimation, demonstrating the robustness of our method.
Further, we analyzed the decay rate of the max-norm error $\|\widetilde{\mathbb{PMI}}-\alpha_{p}\mathbf{V}\mathbf{V}^{\top}\|_{\max}$ for AR(1) by taking the logarithm of this quantity along with the sample size $n$ and sequence length $T$. A linear model was fitted to the median values shown in Figure 3. The results confirm that $\|\widetilde{\mathbb{PMI}}-\alpha_{p}\mathbf{V}\mathbf{V}^{\top}\|_{\max}$ is approximately proportional to $(nT)^{-1/2}\rightarrow 0$, which is consistent with the theoretical results in Theorem 3.6.
Figure 2: The max-norm error between PMI estimators and EHR entity embeddings. In the first row, we fix $T=1000$ and change $n$. In the second row, we fix $n=1000$ and change $T$. Each column represents a distribution of the discourse vectors used in the generative model (2.1).
Figure 3: The data points in this figure represent the median value of the max-norm difference between the low-rank PMI estimator and the population PMI matrix, as shown in the first column of Figure 2.
@@ 4.2 Numerical Results on Hypothesis Testing
In this section, we simulate data following the global null distribution given in Section 2.3.2. For simplicity, we assume that each feature occurs with equal probability, such that $p_{w}=1/d$ for $1\leq w\leq d$. The sets $\{w_{i,t}\}_{1\leq i\leq n,1\leq t\leq T}$ are independently and identically drawn from a $\text{Multinomial}(1,(1/d,\cdots,1/d))$ distribution, considering $d=100$ and $n\in\{400,800,1600\}.$ Under these conditions, it is not reasonable to define the embeddings, and thus our analysis is confined to the empirical estimator. We apply the methods described in Sections 2.3.1 and 2.3.2 to estimate the variance of each entry in $\widehat{\mathbb{PMI}}$ using the Monte-Carlo approach. This involves generating $n$ i.i.d longitudinal code sequences $100$ times, and subsequently calculating the family-wise error rate (FWER) for each method. When patient-level data becomes accessible, we estimate the variance of each entry using (2.9). In contrast, when patient-level data is unavailable, we estimate the variance of each entry via (2.11).
The testing is conducted on all unique edges $(w,w^{\prime}),1\leq w\neq w^{\prime}\leq d$, with the nominal significance level set as $\alpha=0.05$ for each edge. We compare methods with and without using patient-level data for variance estimation and use Bonferroni correction to control the FWER. The FWER results are presented in Figure 4 (Left). We can see that the FWER goes to $\alpha$ as the sample size increases.
Figure 4: The left plot compares the FWER of our proposed method with and without patient-level data under the global null hypothesis; the remaining three plots analyzes the relation between power and signal strength of the proposed method under the alternative hypothesis.
@@ 4.3 Numerical Results on the Asymptotic Normality
In this section, we generate data from (2.1). The embeddings are constructed using $\mathbf{V}=d^{-\kappa}\mathbf{U}$, where $\kappa\in\{1,2\}$ and $\mathbf{U}$ is formed from the orthonormal bases of a $d\times p$ matrix with each element independently drawn from a standard normal distribution. Notably, in this context, we do not utilize $\mathbf{V}\mathbf{V}^{\top}$, and there is no distinction between using $\mathbf{V}$ and $\mathbf{V}-\boldsymbol{1}\mu^{\top}$ for data generation. Consequently, the 22nd and 33rd steps outlined in Section 4.1 are unnecessary for embedding construction.
@@
We further study the relation between the statistical power of our proposed methods and the signal strength $(d^{-\kappa})$. Figure 4 shows that as the signal strength grows, the power emerges from near zero to one.
@@ 5 Applications to Electronic Health Records
We validate the efficacy of our algorithm using EHR data from the Veterans Affairs (VA) Healthcare System consisting of $12.6$ million patients who had at least one visit between $2000$ and $2019$. Patient information in the EHR data is structured as triplets (patient, date, clinical entity), which can be interpreted as sequences of EHR entities. The dataset includes four specific domains of codified information: PheCodes for diseases, Current Procedural Terminology (CPT) codes for procedures, lab test codes, and RxNorm codes for medication prescriptions. Codes occurring fewer than $5000$ times were excluded to minimize noise from infrequent codes, resulting in a final count of $1776$ PheCodes, $224$ CCS codes, $6025$ lab codes (including 382 LOINC codes and 5643 Local lab codes), and $1561$ RxNorm codes as detailed in Hong et al. (2021). Consequently, the dataset contains $d=9586$ EHR entities of interest. These codes were then utilized to construct a co-occurrence matrix, adopting a window size, $q$, of $30$ days following the analysis of Hong et al. (2021).
@@
From the results presented in Table 3, it is evident that our algorithm demonstrates a high power in detecting known-relation pairs. In particular, for several pair categories, such as PheCode-PheCode pairs from PheCode Hierarchy, and related disease-disease pairs, our algorithm achieves an exceptional power of over $85\%$ while all other methods have power lower than $75\%$. Furthermore, our method consistently outperforms the selection based on empirical PMI or cosine-similarity from other benchmarks for all pair categories. This underscores the capability of our algorithm to not only establish a threshold for selecting significant edges but also to leverage the variance information of the estimator, surpassing the performance of methods reliant solely on cosine similarity. Similar patterns can be found in Table S5 in Supplementary S3 with target FDR set as $0.1$.
@@
Figure 6: The estimated low-rank PMI of the top $5,000$ entities with the smallest $p$-values when quantifying their relationships with the Alzheimer’s Disease (PheCode:290.11). The x-axis exhibits the rank of the $p$-values, the blue bars exhibit the estimated low-rank PMI and the red line exhibits $\log(-\log(p$-$\text{value}))$. The codes located to the left of the black line are identified as significantly related to the target feature, AD, at a target FDR of $0.01$.
@@
To validate the accuracy of the KG, we conduct a manual review of the top twenty entities with the smallest $p$-values. This group includes eight PheCodes (diseases) and twelve RXNORMs (drugs). These entities are enumerated in Table S6 in Supplementary S3. A substantial body of literature (Sosa-Ortiz et al., 2012; Bathgate et al., 2001; Nussbaum and Ellis, 2003; Reisberg et al., 2003; Tsuno, 2009; Olin and Schneider, 2002; Xu et al., 2008; Rabinowitz et al., 2007) confirms that these entities have a strong association with AD, thereby validating the precision of our KG.
To further assess the effectiveness of the proposed algorithm in identifying entities associated with the given disease AD, we incorporate a comparison with advanced language models ChatGPT (OpenAI, 2023a) and GPT-4 (OpenAI, 2023b). Specifically, we first select the top $50$ entities related to AD based on different methods, and then randomly sample an equal number of entities from the remaining set as negative samples. We then prompt ChatGPT and GPT-4 to rate the clinical relatedness between these $100$ entities and the target disease AD, outputting a score on a scale of 0 to $1$ for each entity. The evaluation involves computing the Spearman rank correlation between the negative $p$-values generated by our algorithm and the scores from ChatGPT/GPT4. Additionally, we examine correlations between cosine similarities obtained from alternative benchmarks and the ChatGPT/GPT scores, as detailed in Table 4. The results indicate that the rank correlation coefficients for KNIT with ChatGPT and GPT-4 are $0.553$ and $0.582$, respectively, outperforming other methods which all showed correlations less than $0.45$ with the language models. For reference, the rank correlation between ChatGPT and GPT-4 themselves is $0.610$.
@@
Figure 7: Visualization of the generated embedding via t-SNE. The left figure exhibits the first dimension and the second dimension; while the right figure exhibits the second dimension and the third dimension. The points are colored the group of the codified codes.
To better understand the generated embedding, we apply t-SNE (Van der Maaten and Hinton, 2008) to further reduce the embedding dimension and visualize the embeddings. From Figure 7, we can see that the generated embedding does help cluster codified codes into different types of codes. The visualization demonstrates that features with closer embedding distances exhibit higher correlations and stronger relationships, suggesting meaningful topological preservation in the latent space. These findings highlight the practical value of the learned embeddings for various downstream applications, including feature clustering analysis and predictive modeling for disease outcomes.